OpenStudy (anonymous):

What is the 6th term of the geometric sequence where a1 = 1,024 and a4 = –16?

OpenStudy (callisto):

x r x r x r a1 ---> a2 --->a3--->a4 So you can see if you want to get a4 from a1, you need to multiply a1 with r for 3 times So a4 = a1 (r)^3 -16 = 1024 (r)^3 -1/64 = r^3 r = -1/4 as you get the common ratio r, you can apply the general formula to find the 6th term T(n) = ar^(n-1) = 1024 (-1/4)^(6-1) = 1024*(-1/4)^5 = 1024*(-1/1024) = work it out yourself :)

OpenStudy (callisto):

@lgbasallote sorry but i think the general term you've written is wrong...

OpenStudy (lgbasallote):

oh yeah! geometric!!! so sorry...i am so screwed up @_@ i got confused hahaha lemme repeat that

OpenStudy (lgbasallote):

a6 = a1 r^(6-1) first let us find r... (((a1)r)r)r) = a4 ar^3 = -16 1024 r^3 = -16 r^3 = -16/1024 r^3 = -1/64 r = -1/4 a6 = 1024(-1/4)^5 a6 = 1024(-1/1024) a6 = -1 better?

OpenStudy (callisto):

sure :)

OpenStudy (anonymous):

thank you so much for your help!!! means a lot!

OpenStudy (callisto):

(if you are talking to me) : You're welcome! Glad to help :D